OpenStudy (anonymous):

a parabola opens downward and has a vertex at (3,4). the parabola is increasing over the interval.

OpenStudy (anonymous):

which interval?

OpenStudy (anonymous):

answers are: A.all x from -\[\infty \] to 3. B. all x from -\[\infty\]to 4. C. all x from 3 to \[\infty\]. D. all x from 4 to \[\infty\].

OpenStudy (anonymous):

 clearly from -infinity to 3.